Convert the following rectangular coordinates into polar coordinates (r,θ) so that r < 0


and 0 ≤ θ ≤ 2π :


(4,−4√3)




1
Expert's answer
2022-05-08T11:49:48-0400
"r^2=(4)^2+(-4\sqrt{3})^2=64"

"r=-8"

"\theta=\tan^{-1}(\dfrac{-4\sqrt{3}}{4})+\pi=2\pi/3"

"(-8, 2\pi/3)"